OpenStudy (xximhisgirlxx):

Could anyone draw a scatter plot of this table for me? I'll give a medal. Table is in comments.

OpenStudy (xximhisgirlxx):



OpenStudy (xximhisgirlxx):

@Mertsj

jimthompson5910 (jim_thompson5910):

Ok give me a few moments

OpenStudy (xximhisgirlxx):

Okay thank you so much..

jimthompson5910 (jim_thompson5910):

Normally I would use excel, but geogebra does a much better job at scatter plots

OpenStudy (xximhisgirlxx):

Yeah lol

jimthompson5910 (jim_thompson5910):

have you heard of geogebra? or used it?

OpenStudy (xximhisgirlxx):

No I have not

jimthompson5910 (jim_thompson5910):

alright, nearly done

jimthompson5910 (jim_thompson5910):

I was just curious, you should check it out when you have time

OpenStudy (xximhisgirlxx):

Again thank you!

jimthompson5910 (jim_thompson5910):

see attached pic

OpenStudy (xximhisgirlxx):

Could you help me with another problem?

OpenStudy (xximhisgirlxx):

One question is, "What are two equations that model this table"

OpenStudy (xximhisgirlxx):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so you want to find the linear regression equation?

OpenStudy (xximhisgirlxx):

It's just asking What are two equations for this table?

jimthompson5910 (jim_thompson5910):

hmm well off the top of my head, it looks like no simple equation goes through all of the points

jimthompson5910 (jim_thompson5910):

but you can use regression to get a line or curve that's close as possible to each point

jimthompson5910 (jim_thompson5910):

unfortunately regression is a very long and complicated process but luckily geogebra also has regression functions built in, so that saves us a ton of time

jimthompson5910 (jim_thompson5910):

here is a picture of the linear regression line (in green) and the exponential regression curve (in blue)

jimthompson5910 (jim_thompson5910):

the equation of the green linear regression line is y = 1.62121x - 4.73333 this is an approximation of course

jimthompson5910 (jim_thompson5910):

the equation of the blue exponential curve is y = 2.11822e^(0.13693x)

OpenStudy (xximhisgirlxx):

Now it says what are the domain and range of the equations?

jimthompson5910 (jim_thompson5910):

well it looks like each point is in Q1, so that would give off the idea that the domain is the set of positive numbers

jimthompson5910 (jim_thompson5910):

similarly, the range would also be the set of positive numbers (since each point is in Q1)

OpenStudy (xximhisgirlxx):

Could you help me out with two more? :)

OpenStudy (xximhisgirlxx):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (xximhisgirlxx):

About how much money did Stacies store earn in the fifth month?

OpenStudy (xximhisgirlxx):

and explain

jimthompson5910 (jim_thompson5910):

this is based off the same scatter plot right?

OpenStudy (xximhisgirlxx):

yes lol

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (xximhisgirlxx):

ok

jimthompson5910 (jim_thompson5910):

ok so there's no data point for the fifth month otherwise, it would be as simple as looking at the table

jimthompson5910 (jim_thompson5910):

but we have data values for months 4 and 6

jimthompson5910 (jim_thompson5910):

month 4: 5 month 6: 4 this is what the table tells us

jimthompson5910 (jim_thompson5910):

average the two to get (5+4)/2 = 9/2 = 4.5

jimthompson5910 (jim_thompson5910):

so one possible value for month 5 is the number 4.5

OpenStudy (xximhisgirlxx):

Okay next question is : How much money do you expect the store to earn in the 24 month?

jimthompson5910 (jim_thompson5910):

so if y is in thousands of dollars, then she earns 4.5 thousand dollars if y is in millions, then she earns 4.5 thousand dollars

jimthompson5910 (jim_thompson5910):

I meant if y is in millions, then she earns 4.5 million dollars

jimthompson5910 (jim_thompson5910):

plug in x = 24 into each equation I gave you and tell me what you get

OpenStudy (xximhisgirlxx):

No solution ??

OpenStudy (xximhisgirlxx):

I don't know what to do lol, i'm confused ;/

OpenStudy (xximhisgirlxx):

next question is : How much money do you expect the store to earn in the 24 month?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (xximhisgirlxx):

alright

jimthompson5910 (jim_thompson5910):

using the linear regression line, we get y = 1.62121x - 4.73333 y = 1.62121(24) - 4.73333 y = 34.17571

jimthompson5910 (jim_thompson5910):

use the exponential regression equation to get y = 2.11822e^(0.13693x) y = 2.11822e^(0.136939(24)) y = 56.662471504214

jimthompson5910 (jim_thompson5910):

so the projection is either 34.17571 or 56.662471504214 for 24 months (depending on which model you use)

OpenStudy (xximhisgirlxx):

I have to go but thank you so much! How do you give medals?

jimthompson5910 (jim_thompson5910):

just click best response, and yw